40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 29 19:41:38 2008, Yoichi, Update, PSL, FSS loop transfer functions 7x
    Reply  Tue Jul 29 23:04:34 2008, Yoichi, Update, PSL, FSS loop transfer functions OpltfPZTOnlyRaw.epsOpltfPZTOnly.eps
Message ID: 761     Entry time: Tue Jul 29 23:04:34 2008     In reply to: 758
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: FSS loop transfer functions 

Quote:

The measurement of the PZT open loop TF is very suspicious. According to this, the PC path has a very large gain even at very low frequencies (there is no cross over above 1kHz). This cannot be true. Maybe the cavity's optical gain was low when it was locked with only the PZT. I will re-measure it.


I measured it again and found that the loop was oscillating at 13.5kHz. I think this oscillation prevented the ref. cavity from building up the power and consequently lowered the optical gain making it marginally stable. So the PZT path open loop TF posted in the previous entry is wrong.

I was able to stop the oscillation by lowering the gain down to CG=-7.6dB and FG=-8.78dB.
The first attachment shows the measured open loop transfer function.
Since the gain setting is different from when the over all open loop TF was measured, I scaled the gain (attachment 2).
However, this plot seems to have too much gain. Scaling it down by 20dB makes it overlap with the over all open loop TF.
Maybe the gain reading on the EPICS screen is wrong. I will measure the actual gain tomorrow.
Attachment 1: OpltfPZTOnlyRaw.eps  23 kB  | Hide | Hide all
OpltfPZTOnlyRaw.eps
Attachment 2: OpltfPZTOnly.eps  23 kB  | Hide | Hide all
OpltfPZTOnly.eps
ELOG V3.1.3-